# UniDepthV2: Universal Monocular Metric Depth Estimation Made Simpler

### v1

@@ Abstract
Index Terms:
Depth estimation, 3D estimation, camera prediction, geometric perception, foundation model.
@@ II Related Work
Metric and Scale-Agnostic Depth Estimation.  It is crucial to distinguish Monocular Metric Depth Estimation (MMDE) from scale-agnostic, namely up-to-a-scale, monocular depth estimation. MMDE SotA approaches typically confine training and testing to the same domain. However, challenges arise, such as overfitting to the training scenario leading to considerable performance drops in the presence of minor domain gaps, often overlooked in benchmarks like NYU-Depthv2 [22] (NYU) and KITTI [23]. On the other hand, scale-agnostic depth methods, pioneered by MiDaS [24], OmniData [25], and LeReS [26], show robust generalization by training on extensive datasets. The paradigm has been elevated to another level by repurposing depth-conditioned generative methods for RGB to RGB-conditioned depth generative methods [27] or large-scale semi-supervised pre-training as in the DepthAnything series [28, 29]. The limitation of all these methods lies in the absence of a metric output, hindering practical usage in downstream applications.
Monocular Metric Depth Estimation.  The introduction of end-to-end trainable neural networks in MMDE, pioneered by [6], marked a significant milestone, also introducing the optimization process through the Scale-Invariant log loss ($\mathrm{SI}_{\log}$). Subsequent developments witnessed the emergence of advanced networks, ranging from convolution-based architectures [7, 30, 31, 10] to transformer-based approaches [32, 8, 11, 12]. Despite impressive achievements on established benchmarks, MMDE models face challenges in zero-shot scenarios, revealing the need for robust generalization against appearance and geometry domain shifts.
General Monocular Metric Depth Estimation.  Recent efforts focus on developing MMDE models [33, 15, 14] for general depth prediction across diverse domains. These models often leverage camera awareness, either by directly incorporating external camera parameters into computations [34, 15] or by normalizing the shape or output depth based on intrinsic properties, as seen in [35, 36, 14, 16]. A new paradigm recently emerged [17, 37], where the goal is to directly estimate the 3D scene from the input image without any additional information other than the RGB input. Our approach fits in the latter new paradigm, namely universal MMDE: we do not require any additional prior information at test time, such as access to camera information.
@@ III UniDepthV2
The general-purpose nature of our MMDE method requires inferring both depth and camera intrinsics to make 3D predictions based only on imagery observations. We design the 3D output space presenting a natural disentanglement of the two sub-tasks, namely depth estimation and camera calibration. In particular, we exploit the pseudo-spherical representation where the basis is defined by azimuth, elevation, and log-depth, i.e. ($\theta$,$\phi$,$z_{\log}$), in contrast to the Cartesian representation ($x$,$y$,$z$). The strength of the proposed pseudo-spherical representation lies in the decoupling of camera ($\theta$,$\phi$) and depth ($z_{\log}$) components, ensuring their orthogonality by design, in contrast to the entanglement present in Cartesian representation.
@@ III-D Edge-Guided Normalized Loss
Modern depth estimation methods must balance global scene understanding with local geometric precision. While UniDepth excels at the former, it lacks accuracy in local, fine-grained details of the geometry of the depicted scenes. To address this, UniDepthV2 involves a novel loss function, named Edge-Guided Scale-Shift Invariant Loss ($\mathcal{L}_{\mathrm{EG-SSI}}$), which is explicitly designed to enhance local precision. This loss is computed over image patches extracted from regions where the RGB spatial gradient ranks in the top 5%-quantile, capturing high-contrast areas likely to contain depth discontinuities. Patch sizes are randomly sampled between 4% and 8% of the input image’s smallest dimension. By concentrating on these visually salient regions, our model learns to distinguish between genuine geometric discontinuities and misleading high-frequency textures that do not correspond to actual depth changes. For instance, structured patterns such as checkerboard textures or repetitive details on flat surfaces can falsely suggest depth variations, leading to hallucinated discontinuities.
@@
$\mathcal{L}_{\mathrm{EG-SSI}}(\mathbf{D},\mathbf{D^{*}},\Omega)=\sum_{\omega\in\Omega}\left||\mathcal{N}_{\omega}(\mathbf{D}_{\omega})-\mathcal{N}_{\omega}(\mathbf{D}^{*}_{\omega})\right||_{1},$ (3)
@@ III-E Network Design
Architecture.  Our network, described in Fig. 2, comprises an Encoder Backbone, a Camera Module, and a Depth Module. The encoder is ViT-based [38], producing features at four different “scales”, i.e$\{\mathbf{F}_{i}\}^{4}_{i=1}$, with $\mathbf{F}_{i}\in\mathbb{R}^{h\times w\times C}$, where $(h,w)=(\frac{H}{14},\frac{W}{14})$.
@@
The Depth Module is fed with the four feature maps $\{\mathbf{F}_{i}\}^{4}_{i=1}$ from the encoder. Each feature map $\mathbf{F}_{i}$ is conditioned on the camera prompts $\mathbf{E}$ to obtain $\mathbf{D|E}$ as described in Sec. III-B with a different cross-attention layer. The four feature maps are then processed with an FPN-style decoder where the “lateral” convolution is transposed convolution to match the ViT resolution to the resolution of the different layers of the FPN. The log-depth prediction $\mathbf{Z}_{\log}\in\mathbb{R}^{H\times W\times 1}$ corresponds to the last FPN feature map which is upsampled to the original input shape and processed with two convolutional layers. The final 3D output $\mathbf{O}\in\mathbb{R}^{H\times W\times 3}$ is the concatenation of predicted rays and depth, $\mathbf{O}=\mathbf{C}||\mathbf{Z}$, with $\mathbf{Z}$ as element-wise exponentiation of $\mathbf{Z}_{\log}$.
Optimization.  The optimization process is guided by a re-formulation of the Mean Squared Error (MSE) loss in the final 3D output space ($\theta$,$\phi$,$z_{\log}$) from Sec. III-A as:
$\vspace{-4pt}\begin{split}\mathcal{L}_{\lambda\mathrm{MSE}}(\bm{\varepsilon})=\|\mathbb{V}[\bm{\varepsilon}]\|_{1}+\bm{\lambda}^{T}(\mathbb{E}[\bm{\varepsilon}]\odot\mathbb{E}[\bm{\varepsilon}]),\end{split}$ (4)
where $\bm{\varepsilon}=\hat{\mathbf{o}}-\mathbf{o}^{*}\in\mathbb{R}^{3}$, $\hat{\mathbf{o}}=(\hat{\theta},\hat{\phi},\hat{z}_{\log})$ is the predicted 3D output, $\mathbf{o}^{*}=(\theta^{*},\phi^{*},z_{\log}^{*})$ is the GT 3D value, and $\bm{\lambda}=(\lambda_{\theta},\lambda_{\phi},\lambda_{z})\in\mathbb{R}^{3}$ is a vector of weights for each dimension of the output. $\mathbb{V}[\bm{\varepsilon}]$ and $\mathbb{E}[\bm{\varepsilon}]$ are computed as the vectors of empirical variances and means for each of the three output dimensions over all pixels, i.e$\{\bm{\varepsilon}^{i}\}_{i=1}^{N}$. Note that if $\lambda_{d}=1$ for a dimension $d$, the loss represents the standard MSE loss for that dimension. If $\lambda_{d}<1$, a scale-invariant loss term is added to that dimension if it is expressed in log space, e.g. for the depth dimension $z_{\log}$, or a shift-invariant loss term is added if that output is expressed in linear space. In particular, if only the last output dimension is considered, i.e. the one corresponding to depth, and $\lambda_{z}=0.15$ is utilized, the corresponding loss is the standard $\mathrm{SI}_{\log}$. In our experiments, we set $\lambda_{\theta}=\lambda_{\phi}=1$ and $\lambda_{z}=0.15$. In addition, we extended the optimization with the supervision for the uncertainty prediction, defined as an L1 loss between the predicted uncertainty and the detached error in log space between predicted depth ($\mathbf{Z}_{\log}$) and GT depth ($\mathbf{Z}^{*}_{\log}$). More formally,
@@ IV-A Experimental Setup
Data.  The training data is the combination of 24 publicly available datasets: A2D2 [39], Argoverse2 [40], ARKit-Scenes [41], BEDLAM [42], BlendedMVS [43], DL3DV [44], DrivingStereo [45], DynamicReplica [46], EDEN [47], HOI4D [48], HM3D [49], Matterport3D [50], Mapillary-PSD [36], MatrixCity [51], MegaDepth [52], NianticMapFree [53], PointOdyssey [54], ScanNet [55], ScanNet++ [56], TartanAir [57], Taskonomy [58], Waymo [59], and WildRGBD [60] for a total of 16M images. We evaluate the generalizability of models by testing them on 8 datasets not seen during training, grouped in different domains that are defined based on indoor or outdoor settings. The indoor group corresponds to the validation splits of SUN-RGBD [61], IBims [62], TUM-RGBD [63], and HAMMER [64], while the outdoor group comprises ETH3D [65], Sintel [66], DDAD [67], and NuScenes [68].
@@
Evaluation Details.  All methods have been re-evaluated with a fair and consistent pipeline. In particular, we do not exploit any test-time augmentations and we utilize the same weights for all zero-shot evaluations. We use the checkpoint corresponding to the zero-shot model for each method, i.e. not fine-tuned on KITTI or NYU. The metrics utilized in the main experiments are $\mathrm{\delta_{1}^{SSI}}$, $\mathrm{F_{A}}$, and $\mathrm{\rho_{A}}$. $\mathrm{\delta_{1}}$ measures the depth estimation performance. $\mathrm{F_{A}}$ is the area under the curve (AUC) of F1-score [69] up to $1/20$ of the datasets’ maximum depth and evaluates 3D estimation accuracy. $\mathrm{\rho_{A}}$ evaluates the camera performance and is the AUC of the average angular error of camera rays up to 15. We do not use parametric evaluation of e.g. focal length, since it is a less flexible metric across diverse camera models and perfectly unrectified images. Moreover, we present the fine-tuning ability of UniDepthV2 by training the final checkpoint on KITTI and NYU-Depth V2 and evaluating in-domain, as per standard practice.
@@
Implementation Details.  UniDepthV2 is implemented in PyTorch [70] and CUDA [71]. For training, we use the AdamW [72] optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$) with an initial learning rate of $5\times{}10^{-5}$. The learning rate is divided by a factor of 10 for the backbone weights for every experiment and weight decay is set to $0.1$. We exploit Cosine Annealing as learning rate and weight decay scheduler to one-tenth starting from 30% of the whole training. We run 300k optimization iterations with a batch size of 128. The training time amounts to 6 days on 16 NVIDIA 4090 with half precision. The dataset sampling procedure follows a weighted sampler, where the weight of each dataset is its number of scenes. Our augmentations are both geometric and photometric, i.e. random resizing, cropping, and translation for the former type, and brightness, gamma, saturation, and hue shift for the latter. We randomly sample the image ratio per batch between 2:1 and 1:2. Our ViT [38] backbone is initialized with weights from DINO-pre-trained [73] models. For the ablations, we run 100k training steps with a ViT-S backbone, with the same training pipeline as for the main experiments.

### v2

@@ II Related Work
Metric and Scale-Agnostic Depth Estimation.  It is crucial to distinguish Monocular Metric Depth Estimation (MMDE) from scale-agnostic, namely up-to-a-scale, monocular depth estimation. MMDE SotA approaches typically confine training and testing to the same domain. However, challenges arise, such as overfitting to the training scenario, leading to considerable performance drops in the presence of minor domain gaps, often overlooked in benchmarks like NYU-Depthv2 [22] (NYU) and KITTI [23]. On the other hand, scale-agnostic depth methods, pioneered by MiDaS [24], OmniData [25], and LeReS [26], show robust generalization by training on extensive datasets. The paradigm has been elevated to another level by repurposing depth-conditioned generative methods for RGB to RGB-conditioned depth generative methods [27] or large-scale semi-supervised pre-training as in the DepthAnything series [28, 29]. In particular, these two paradigms have been used extensively for other dense prediction tasks, such as segmentation [30] and surface normal estimation [31, 32], and for other modalities, notably video [33, 34]. The limitation of all these methods lies in the absence of a metric output, hindering practical usage in downstream applications.
Monocular Metric Depth Estimation.  The introduction of end-to-end trainable neural networks in MMDE, pioneered by [6], marked a significant milestone, also introducing the optimization process through the Scale-Invariant log loss ($\mathrm{SI}_{\log}$). Subsequent developments witnessed the emergence of advanced networks, ranging from convolution-based architectures [7, 35, 36, 10] to transformer-based approaches [37, 8, 11, 12, 38]. Despite impressive achievements on established benchmarks, MMDE models face challenges in zero-shot scenarios, revealing the need for robust generalization against appearance and geometry domain shifts.
General Monocular Metric Depth Estimation.  Recent efforts focus on developing MMDE models [39, 15, 14] for general depth prediction across diverse domains. These models often leverage camera awareness, either by directly incorporating external camera parameters into computations [40, 15] or by normalizing the shape or output depth based on intrinsic properties, as seen in [41, 42, 14, 16]. A new paradigm recently emerged [17, 43, 44], where the goal is to directly estimate the 3D scene from the input image without any additional information other than the RGB input. Our approach fits in the latter new paradigm, namely universal MMDE: we do not require any additional prior information at test time, such as access to camera information.
Depth Estimation under Challenging Conditions.  A parallel line of work targets robustness when image formation departs from the Lambertian, well-lit assumption. Diffusion-augmented training and guidance improve performance in rare or adverse conditions (e.g. night, fog, rain, low light) [45], while dedicated training/evaluation protocols formalize the “challenging conditions” setting and report consistent gains across benchmarks [46]. For non-Lambertian materials, methods tailored to transparent and mirror surfaces mitigate severe failure modes [47, 48]. Our work is complementary: we pursue universal MMDE, aiming for a single camera-free model that generalizes across such conditions through large-scale training.
@@ III UniDepthV2
The general-purpose nature of our MMDE method requires inferring both depth and camera intrinsics to make 3D predictions based only on imagery observations. We design the 3D output space presenting a natural disentanglement of the two sub-tasks, namely depth estimation and camera calibration. In particular, we exploit the pseudo-spherical representation where the basis is defined by azimuth, elevation, and log-depth, i.e. ($\theta$$\phi$$z_{\log}$), in contrast to the Cartesian representation ($x$,$y$,$z$). The strength of the proposed pseudo-spherical representation lies in the decoupling of camera ($\theta$$\phi$) and depth ($z_{\log}$) components, ensuring their orthogonality by design, in contrast to the entanglement present in Cartesian representation.
@@ III-D Edge-Guided Normalized Loss
Modern depth estimation methods must balance global scene understanding with local geometric precision. While UniDepth excels at the former, it lacks accuracy in local, fine-grained details of the geometry of the depicted scenes. To address this, UniDepthV2 involves a novel loss function, named Edge-Guided Scale-Shift Invariant Loss ($\mathcal{L}_{\mathrm{EG-SSI}}$), which is explicitly designed to enhance local precision. This loss is computed over image patches extracted from regions where the RGB spatial gradient ranks in the top 5%-quantile, capturing high-contrast areas likely to contain depth discontinuities. Patch sizes are randomly sampled between 4% and 8% of the input image’s smallest dimension. By concentrating on these visually salient regions, our model learns to distinguish between genuine geometric discontinuities and misleading high-frequency textures that do not correspond to actual depth changes. The loss oversamples high RGB-gradient regions so that thin, high-frequency boundaries are well represented. We do not use superpixels (e.g. SLIC [49]) because their emphasis on region uniformity suppresses thin structures and creates a non-maximum-suppression-like effect that harms boundary detail. For instance, structured patterns such as checkerboard textures or repetitive details on flat surfaces can falsely suggest depth variations, leading to hallucinated discontinuities.
@@
[…]
@@ III-E Network Design
Architecture.  Our network, described in Fig. 2, comprises an Encoder Backbone, a Camera Module, and a Depth Module. The encoder is ViT-based [50], producing features at four different “scales”, i.e. $\{\mathbf{F}_{i}\}^{4}_{i=1}$, with $\mathbf{F}_{i}\in\mathbb{R}^{h\times w\times C}$, where $(h,w)=(\frac{H}{14},\frac{W}{14})$.
@@
The Depth Module is fed with the four feature maps $\{\mathbf{F}_{i}\}^{4}_{i=1}$ from the encoder. Each feature map $\mathbf{F}_{i}$ is conditioned on the camera prompts $\mathbf{E}$ to obtain $\mathbf{D|E}$ as described in Sec. III-B with a different cross-attention layer. The four feature maps are then processed with an FPN-style decoder where the “lateral” convolution is transposed convolution to match the ViT resolution to the resolution of the different layers of the FPN. The log-depth prediction $\mathbf{Z}_{\log}\in\mathbb{R}^{H\times W\times 1}$ corresponds to the last FPN feature map which is upsampled to the original input shape and processed with two convolutional layers. The final 3D output $\mathbf{O}\in\mathbb{R}^{H\times W\times 3}$ is the concatenation of predicted rays and depth, $\mathbf{O}=\mathbf{C}||\mathbf{Z}$, with $\mathbf{Z}$ as element-wise exponentiation of $\mathbf{Z}_{\log}$. Owing to the architecture’s modularity, the ray bundle $\mathbf{C}$ need not come from UniDepthV2’s Camera Module: at inference, it can be injected from any camera model with known parameters and a specified unprojection operator.
Optimization.  The optimization process is guided by a re-formulation of the Mean Squared Error (MSE) loss in the final 3D output space ($\theta$$\phi$$z_{\log}$) from Sec. III-A as:
$\vskip-4.0pt\begin{split}\mathcal{L}_{\lambda\mathrm{MSE}}(\bm{\varepsilon})=\|\mathbb{V}[\bm{\varepsilon}]\|_{1}+\bm{\lambda}^{T}(\mathbb{E}[\bm{\varepsilon}]\odot\mathbb{E}[\bm{\varepsilon}]),\end{split}$ (4)
where $\bm{\varepsilon}=\hat{\mathbf{o}}-\mathbf{o}^{*}\in\mathbb{R}^{3}$, $\hat{\mathbf{o}}=(\hat{\theta},\hat{\phi},\hat{z}_{\log})$ is the predicted 3D output, $\mathbf{o}^{*}=(\theta^{*},\phi^{*},z_{\log}^{*})$ is the GT 3D value, and $\bm{\lambda}=(\lambda_{\theta},\lambda_{\phi},\lambda_{z})\in\mathbb{R}^{3}$ is a vector of weights for each dimension of the output. $\mathbb{V}[\bm{\varepsilon}]$ and $\mathbb{E}[\bm{\varepsilon}]$ are computed as the vectors of empirical variances and means for each of the three output dimensions over all pixels, i.e. $\{\bm{\varepsilon}^{i}\}_{i=1}^{N}$. Note that if $\lambda_{d}=1$ for a dimension $d$, the loss represents the standard MSE loss for that dimension. If $\lambda_{d}<1$, a scale-invariant loss term is added to that dimension if it is expressed in log space, e.g. for the depth dimension $z_{\log}$, or a shift-invariant loss term is added if that output is expressed in linear space. In particular, if only the last output dimension is considered, i.e. the one corresponding to depth, and $\lambda_{z}=0.15$ is utilized, the corresponding loss is the standard $\mathrm{SI}_{\log}$. In our experiments, we set $\lambda_{\theta}=\lambda_{\phi}=1$ and $\lambda_{z}=0.15$. In addition, we extended the optimization with the supervision for the uncertainty prediction $\Sigma$, defined as an L1 loss between the predicted uncertainty and the detached error in log space between predicted depth ($\mathbf{Z}_{\log}$) and GT depth ($\mathbf{Z}^{*}_{\log}$). More formally,
@@ IV Experiments
Method SUNRGBD IBims-1 TUM-RGBD
$\mathrm{\delta_{1}}\uparrow$ $\mathrm{ARel}\downarrow$ $F_{1}\uparrow$ $\rho\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{ARel}\downarrow$ $F_{1}\uparrow$ $\rho\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{ARel}\downarrow$ $F_{1}\uparrow$ $\rho\uparrow$
Metric3D†‡ [14] $1.9$ $48.7$ - - $75.1$ $19.3$ - - $7.7$ $61.1$ - -
Metric3Dv2†‡ [16] $81.2$ $13.3$ - - $68.4$ $20.7$ - - $63.0$ $\mathbf{9.3}$ - -
ZoeDepth [39] $80.9$ $13.6$ - - $49.8$ $21.5$ - - $55.6$ $33.6$ - -
UniDepth [17] $94.3$ $10.4$ $78.6$ $85.8$ $15.7$ $41.0$ $30.3$ $\mathbf{76.6}$ $72.3$ $\underline{17.2}$ $54.8$ $86.8$
MASt3R [51] $80.1$ $14.4$ $71.5$ $\underline{92.0}$ $61.0$ $19.7$ $55.7$ $76.0$ $52.4$ $27.9$ $44.1$ $\underline{93.7}$
DepthPro [43] $83.1$ $13.3$ $71.1$ $89.3$ $82.3$ $17.0$ $62.8$ $75.9$ $56.9$ $19.9$ $48.1$ $\mathbf{96.5}$
UniDepthV2-Small $90.8$ $10.5$ $74.2$ $87.7$ $86.6$ $13.5$ $62.4$ $67.5$ $69.0$ $23.6$ $50.6$ $86.1$
UniDepthV2-Base $\underline{94.4}$ $\underline{8.4}$ $\underline{79.9}$ $91.1$ $\underline{89.7}$ $\underline{11.1}$ $\underline{68.5}$ $\underline{76.5}$ $\underline{77.5}$ $20.7$ $\underline{57.3}$ $89.4$
UniDepthV2-Large $\mathbf{96.4}$ $\mathbf{6.8}$ $\mathbf{84.6}$ $\mathbf{93.4}$ $\mathbf{94.5}$ $\mathbf{7.8}$ $\mathbf{70.9}$ $74.1$ $\mathbf{90.5}$ $22.1$ $\mathbf{62.9}$ $89.6$
TABLE II: Results for Outdoor Domains. All methods are tested in a zero-shot fashion. Missing values (-) indicate the model’s inability to produce the respective output. †: Requires ground-truth (GT) camera for 3D reconstruction. ‡: Requires GT camera for 2D depth map inference. For DDAD, Metric3D and Metric3Dv2 are excluded from evaluation as they are trained on it.
Method ETH3D Sintel DDAD NuScenes
$\mathrm{\delta_{1}}\uparrow$ $\mathrm{ARel}\downarrow$ $\mathrm{F_{1}}\uparrow$ $\mathrm{\rho}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{ARel}\downarrow$ $\mathrm{F_{1}}\uparrow$ $\mathrm{\rho}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{ARel}\downarrow$ $\mathrm{F_{1}}\uparrow$ $\mathrm{\rho}\uparrow$ $\mathrm{\delta_{1}}\uparrow$ $\mathrm{ARel}\downarrow$ $\mathrm{F_{1}}\uparrow$ $\mathrm{\rho}\uparrow$
Metric3D†‡ [14] $19.7$ $136.8$ - - $1.4$ $105.5$ - - n/a n/a n/a n/a $75.4$ $23.7$ - -
Metric3Dv2†‡ [16] $\mathbf{90.0}$ $\mathbf{12.7}$ - - $\mathbf{34.5}$ $\mathbf{43.9}$ - - n/a n/a n/a n/a $84.1$ $23.6$ - -
ZoeDepth [39] $33.8$ $54.7$ - - $5.6$ $95.9$ - - $27.9$ $50.9$ - - $33.8$ $42.0$ - -
UniDepth [17] $18.5$ $53.3$ $27.6$ $42.6$ $13.2$ $124.6$ $40.2$ $65.6$ $\underline{85.8}$ $\mathbf{12.5}$ $\underline{72.8}$ $\mathbf{98.1}$ $84.6$ $\mathbf{12.7}$ $\underline{64.4}$ $\mathbf{97.7}$
MASt3R [51] $21.4$ $45.3$ $28.4$ $\underline{92.2}$ $17.2$ $71.6$ $41.5$ $72.2$ $4.3$ $56.7$ $22.1$ $74.6$ $2.7$ $65.6$ $13.6$ $78.3$
DepthPro [43] $39.7$ $65.2$ $41.2$ $77.4$ $26.2$ $133.4$ $49.7$ $75.2$ $29.9$ $37.3$ $42.1$ $83.0$ $56.6$ $28.7$ $46.5$ $79.1$
UniDepthV2-Small $64.6$ $113.2$ $44.3$ $78.4$ $14.6$ $107.7$ $37.1$ $73.5$ $83.3$ $16.4$ $68.5$ $94.7$ $82.1$ $18.3$ $59.7$ $96.2$
UniDepthV2-Base $75.4$ $70.1$ $\underline{53.5}$ $91.4$ $31.9$ $\underline{50.1}$ $\mathbf{51.8}$ $\underline{75.9}$ $\underline{86.8}$ $14.2$ $71.4$ $96.1$ $\underline{85.3}$ $16.2$ $63.6$ $96.6$
UniDepthV2-Large $\underline{85.2}$ $\underline{16.6}$ $\mathbf{59.3}$ $\mathbf{92.6}$ $\underline{34.4}$ $61.0$ $\underline{51.4}$ $\mathbf{76.3}$ $\mathbf{88.2}$ $\underline{13.8}$ $\mathbf{73.3}$ $\underline{96.7}$ $\mathbf{87.0}$ $\underline{15.0}$ $\mathbf{66.7}$ $\underline{97.2}$
@@ IV-A Experimental Setup
Data.  The training data is the combination of 23 publicly available datasets: A2D2 [52], Argoverse2 [53], ARKit-Scenes [54], BEDLAM [55], BlendedMVS [56], DL3DV [57], DrivingStereo [58], DynamicReplica [59], EDEN [60], HOI4D [61], HM3D [62], Matterport3D [63], Mapillary-PSD [42], MatrixCity [64], MegaDepth [65], NianticMapFree [66], PointOdyssey [67], ScanNet [68], ScanNet++ [69], TartanAir [70], Taskonomy [71], Waymo [72], and WildRGBD [73] for a total of 16M images. We evaluate the generalizability of models by testing them on ten datasets not seen during training, grouped in different domains that are defined based on indoor, outdoor or “challenging” settings. The indoor group corresponds to the validation splits of SUN-RGBD [74], IBims [75], and TUM-RGBD [76], the outdoor group comprises ETH3D [77], Sintel [78], DDAD [79], and NuScenes [80], while the “challenging” domain is composed of HAMMER [81], Booster [48], and FLSea [82].
@@
Evaluation Details.  All methods have been re-evaluated with a fair and consistent pipeline. In particular, we do not exploit any test-time augmentations, and we utilize the same weights for all zero-shot evaluations. We use the checkpoint corresponding to the zero-shot model for each method, i.e. not fine-tuned on KITTI or NYU. The metrics utilized in the main experiments are $\mathrm{\delta_{1}^{SSI}}$, $\mathrm{F_{A}}$, and $\mathrm{\rho_{A}}$. $\mathrm{\delta_{1}}$ measures the depth estimation performance. $\mathrm{F_{A}}$ is the area under the curve (AUC) of F1-score [83] up to $1/20$ of the datasets’ maximum depth and evaluates 3D estimation accuracy. $\mathrm{\rho_{A}}$ [84, 85] evaluates the camera performance and is the AUC of the average angular error of camera rays up to 15. We do not use parametric evaluation of e.g. focal length, since it is a less flexible metric across diverse camera models and perfectly unrectified images. Moreover, we present the fine-tuning ability of UniDepthV2 by training the final checkpoint on KITTI and NYU-Depth V2 and evaluating in-domain, as per standard practice.
Confidence predictions are evaluated with specific and established metrics such as AUSE [86] (w.r.t. $\delta_{1}$) and its normalized version (nAUSE) and Spearman’s $\rho$. $\mathrm{AUSE}$ ranks pixels by predicted uncertainty, progressively masks the most uncertain fraction, recomputes $\delta_{1}$ at each step, and integrates the gap to the oracle curve (ranking by true error). Oracle’s AUSE is 0 by definition and $\mathrm{nAUSE}$ is $\mathrm{AUSE}$ normalized by the random error, i.e. $\mathrm{nAUSE}=1$ equals random. $\rho$ is the rank correlation between predicted uncertainty and per-pixel error ($1$ perfect ordering, 0 no monotonic relation, <0<0 inverted). Both metrics therefore assess ranking quality: AUSE under progressive masking and $\rho$ as a global monotonicity check.
@@
Implementation Details.  UniDepthV2 is implemented in PyTorch [87] and CUDA [88]. For training, we use the AdamW [89] optimizer ($\beta_{1}=0.9$, $\beta_{2}=0.999$) with an initial learning rate of $5\times{}10^{-5}$. The learning rate is divided by a factor of 10 for the backbone weights for every experiment and weight decay is set to $0.1$. We exploit Cosine Annealing as learning rate and weight decay scheduler to one-tenth starting from 30% of the whole training. We run 300k optimization iterations with a batch size of 128. The training time amounts to 6 days on 16 NVIDIA 4090 with half precision. The dataset sampling procedure follows a weighted sampler, where the weight of each dataset is its number of scenes. Our augmentations are both geometric and photometric, i.e. random resizing, cropping, and translation for the former type, and brightness, gamma, saturation, and hue shift for the latter. We randomly sample the image ratio per batch between 2:1 and 1:2. Our ViT [50] backbone is initialized with weights from DINO-pretrained [90] models. We train three models, one for each available ViT backbone, i.e. Small, Base, and Large, corresponding to the UniDepthV2 variants “-Small”, “-Base”, and “-Large”. For the ablations, we run 100k training steps with a ViT-S backbone, with the same training pipeline as for the main experiments. Because the ablation runs use a shorter schedule and a fixed ViT-S backbone, the ablation “final” model is not directly comparable to the main UniDepthV2-Small results in Tabs. I, II and III.